# Implicit Bias and Invariance: How Hopfield Networks Efficiently Learn Graph Orbits

### v1

@@ 4.3 Invariant parameters
Before proceeding we pause to reflect on the implications of our results with respect to (hillar2021hidden, Conjecture 1). Together, Theorem 3.2, Lemma 4.3 and Lemma 4.4 imply that $k$-cliques on $v$ vertices can be strictly memorized with high probability as long as $N=\tilde{\Omega}(v^{3}k^{2})$. If $k=\alpha v$, where $\alpha\in(0,1)$ is a constant and we assume $\alpha v$ is an integer, then using Stirling’s approximation the critical ratio satisfies $\tilde{O}(v^{5})/\binom{v}{\alpha v}=\tilde{O}\left(2^{-vH(\alpha)}v^{5.5}\right)$ where here $H$ denotes the binary entropy function. Clearly the critical ratio decays to zero at a rate which as $v\to\infty$ is dominated by the exponential term. Our experiments and results thus far suggest that memorization of a graph isomorphism class occurs when the training sample is sufficiently large that the optimizer is forced to return a solution lying close to the invariant set $\Psi(\Phi_{n})\subset F(\mathbb{R}^{3})$. The following lemma establishes that the HSVM solution on the full isomorphism class, which as $N\rightarrow\infty$ is equivalent to the training sample with probability one, must be graph isomorphism invariant.
@@ A.2 Learning algorithms for Hopfield Networks
We briefly describe several classical learning rules [Hertz91] that can be applied to find parameters in Hopfield networks. These methods typically trade off between biological plausibility and performance. We remark that this list is far from exhaustive; see [Tolmachev20] for a recent summary.
  • Outer-Product Rule [hebb1949organization, Nakano72, amari1972learning, hopfield1982neural]. In the attractor neural network case [hopfield1982neural], this Hebbian rule constructs weights as the normalized sum of training pattern outer products. This rule is simple, biologically motivated, and local in nature, but it is often observed to suffer from spurious attractors, shallow basins of attraction, and overall limited capacity.
  • Perceptron Rule [rosenblatt1958perceptron]. The difference between the desired response – that the network dynamics should fix a training sample – and the actual linear-threshold response of a neuron gives a learning signal to update parameters.
  • Delta [hoff1960adaptive, rescorla1972theory]. The delta rule, also called the Mean Squared Error (MSE) or Least Mean Square (LMS) rule, considers a relaxation and follows a gradient to minimize the squared error between the linear output activations of neurons and the training pattern to memorize.
  • Projection Rule [personnaz1985storage, personnaz1986collective]. The weight matrix is obtained by projecting onto the span of the training data and then zeroing the diagonal entries.
  • Storkey Rule [storkey1997increasing, storkey1999novel]. A modification of the Hebbian update that reduces interference between patterns by accounting for previously stored ones. This rule achieves higher storage capacity than Hebbian learning and reduces spurious minima.
@@ A.6 Hoeffding’s inequality in Hilbert space
[Specialization of [pinelis94, Thm. 3.5] For $i\in[N]$ and $R\in\mathbb{R}_{>0}$, let ${\bm{x}}_{i}\in\mathbb{R}^{q}$ be independent, mean zero random vectors which satisfy $\|{\bm{x}}_{i}\|\leq R$ almost surely. Let $S_{N}=\sum_{i=1}^{N}{\bm{x}}_{i}$, then for $t\in\mathbb{R}_{\geq 0}$ we have
@@ Appendix B Proofs of results
Note by construction that $g\in\mathcal{G}_{\Lambda}$ implies $g:\mathbb{R}^{n}\rightarrow[0,1]$. We now compute the empirical Rademacher complexity of $\mathcal{G}_{\Lambda}$ on a sample $\mathcal{S}_{N}=({\bm{x}}_{i})_{i\in[N]}$, to this end let ${\bm{\sigma}}\in\{\pm 1\}^{n}$ and ${\bm{\epsilon}}\in\{\pm\}^{N\times n}$ be a random vector and matrix respectively whose entries are mutually i.i.d. and distributed uniformly on $\{\pm 1\}$. As $\ell$ is 1-Lipschitz in the $\ell_{2}$ norm, then applying a vector contraction inequality [vector-contrac, Corollary 1] we have
@@
Let $\delta\in\mathbb{R}_{>0}$. Applying a Rademacher complexity bound, e.g., [FoML, Thm 3.3], then with probability at least $1-\delta$ over the random sample $\mathcal{S}_{N}$
@@ B.4.2 Proximity of AHSVM solution to the invariant set
where the last inequality follows from Lemma A.7 and the assumption $m\geq 2$. We now deploy Lemma A.9, a specialization of [pinelis94, Th, 3.5]. In particular, given some $\epsilon\in\mathbb{R}_{\geq 0}$ and letting $S_{N}=\sum_{i=1}^{N}(\bar{{\bm{u}}}({\bm{x}})-{\bm{\mu}})$ then
@@ C.3 Clique generalization in DAMs
We conducted experiments comparing the generalization performance of DAMs with that of MEF when presented with increasing numbers of cliques as training data. We used the architecture and algorithm described in [krotov2016dense] for polynomial degree $6$ in the energy function. The results in Fig. 6 show very different generalization behaviors. These preliminary investigations suggest that DAMs do not generalize well in the setting of cliques, but much more work is needed to understand their behavior.

### v2

@@ 4.3 Invariant parameters
Before proceeding we pause to reflect on the implications of our results with respect to (Hillar et al., 2021, Conjecture 1). Together, Theorem 3.2, Lemma 4.3 and Lemma 4.4 imply that $k$-cliques on $v$ vertices can be strictly memorized with high probability as long as $N=\tilde{\Omega}(v^{3}k^{2})$. If $k=\alpha v$, where $\alpha\in(0,1)$ is a constant and we assume $\alpha v$ is an integer, then using Stirling’s approximation the critical ratio satisfies $\tilde{O}(v^{5})/\binom{v}{\alpha v}=\tilde{O}\left(2^{-vH(\alpha)}v^{5.5}\right)$ where here $H$ denotes the binary entropy function. Clearly the critical ratio decays to zero at a rate which as $v\to\infty$ is dominated by the exponential term. Our experiments and results thus far suggest that memorization of a graph isomorphism class occurs when the training sample is sufficiently large that the optimizer is forced to return a solution lying close to the invariant set $\Psi(\Phi_{n})\subset F(\mathbb{R}^{3})$. The following lemma establishes that the HSVM solution on the full isomorphism class, which as $N\rightarrow\infty$ is equivalent to the training sample with probability one, must be graph isomorphism invariant.
@@ A.2 Learning algorithms for Hopfield Networks
We briefly describe several classical learning rules [15] that can be applied to find parameters in Hopfield networks. These methods typically trade off between biological plausibility and performance. We remark that this list is far from exhaustive; see [48] for a recent summary.
  • Outer-Product Rule [14, 34, 1, 20]. In the attractor neural network case [20], this Hebbian rule constructs weights as the normalized sum of training pattern outer products. This rule is simple, biologically motivated, and local in nature, but it is often observed to suffer from spurious attractors, shallow basins of attraction, and overall limited capacity.
  • Perceptron Rule [43]. The difference between the desired response – that the network dynamics should fix a training sample – and the actual linear-threshold response of a neuron gives a learning signal to update parameters.
  • Delta [19, 42]. The delta rule, also called the Mean Squared Error (MSE) or Least Mean Square (LMS) rule, considers a relaxation and follows a gradient to minimize the squared error between the linear output activations of neurons and the training pattern to memorize.
  • Projection Rule [39, 38]. The weight matrix is obtained by projecting onto the span of the training data and then zeroing the diagonal entries.
  • Storkey Rule [47, 46]. A modification of the Hebbian update that reduces interference between patterns by accounting for previously stored ones. This rule achieves higher storage capacity than Hebbian learning and reduces spurious minima.
@@ A.6 Hoeffding’s inequality in Hilbert space
[Specialization of [40, Thm. 3.5] For $i\in[N]$ and $R\in\mathbb{R}_{>0}$, let ${\bm{x}}_{i}\in\mathbb{R}^{q}$ be independent, mean zero random vectors which satisfy $\|{\bm{x}}_{i}\|\leq R$ almost surely. Let $S_{N}=\sum_{i=1}^{N}{\bm{x}}_{i}$, then for $t\in\mathbb{R}_{\geq 0}$ we have
@@ Appendix B Proofs of results
Note by construction that $g\in\mathcal{G}_{\Lambda}$ implies $g:\mathbb{R}^{n}\rightarrow[0,1]$. We now compute the empirical Rademacher complexity of $\mathcal{G}_{\Lambda}$ on a sample $\mathcal{S}_{N}=({\bm{x}}_{i})_{i\in[N]}$, to this end let ${\bm{\sigma}}\in\{\pm 1\}^{n}$ and ${\bm{\epsilon}}\in\{\pm\}^{N\times n}$ be a random vector and matrix respectively whose entries are mutually i.i.d. and distributed uniformly on $\{\pm 1\}$. As $\ell$ is 1-Lipschitz in the $\ell_{2}$ norm, then applying a vector contraction inequality [29, Corollary 1] we have
@@
Let $\delta\in\mathbb{R}_{>0}$. Applying a Rademacher complexity bound, e.g., [32, Thm 3.3], then with probability at least $1-\delta$ over the random sample $\mathcal{S}_{N}$
@@ B.4.2 Proximity of AHSVM solution to the invariant set
where the last inequality follows from Lemma A.7 and the assumption $m\geq 2$. We now deploy Lemma A.9, a specialization of [40, Th, 3.5]. In particular, given some $\epsilon\in\mathbb{R}_{\geq 0}$ and letting $S_{N}=\sum_{i=1}^{N}(\bar{{\bm{u}}}({\bm{x}})-{\bm{\mu}})$ then
@@ C.3 Clique generalization in DAMs
We conducted experiments comparing the generalization performance of DAMs with that of MEF when presented with increasing numbers of cliques as training data. We used the architecture and algorithm described in [25] for polynomial degree $6$ in the energy function. The results in Fig. 6 show very different generalization behaviors. These preliminary investigations suggest that DAMs do not generalize well in the setting of cliques, but much more work is needed to understand their behavior.